# ViSaRL: Visual Reinforcement Learning Guided by Human Saliency

### v1

@@ I INTRODUCTION
A key ingredient in solving visual control tasks is to learn visual representations that capture useful features of the sensory input to simplify the decision making process. Many works in the deep reinforcement learning (RL) community have proposed to learn such representations through various self-supervised objectives including contrastive learning [3] and data augmentation [4]. By contrast, we focus on self-supervision using saliency as additional human domain knowledge to inform the representation of task-relevant features in the visual input while filtering out perceptual noise.
@@
Figure 1: ViSaRL trains a saliency prediction model from a few human-annotated saliency maps. This model is used to augment an offline image dataset with saliency. A visual encoder is pretrained with the paired image and saliency dataset and used during downstream policy learning to generate latent representations of the agent’s observations.
@@ II RELATED WORK
User Interfaces for Human Saliency. ViSaRL needs a small number of human-annotated saliency maps to bootstrap the saliency prediction network (Figure 1). Prior work used superpixel segmentation [22] to first divide each image into segments, and then asked humans to click on the segments that are salient [17]. However, that method requires manually checking and combining the segments that belong to the same object before showing the images to annotators, burdening system designers. As an alternative, Boyd et al. [23, 20] used interfaces where the annotators created binary masks by simply clicking on images. We employ a similar but simpler interface: an annotator clicks on the salient parts of the image, and a Gaussian kernel is applied around selected pixels to achieve smooth saliency maps.
Representation Learning for RL. Saliency maps are essentially representations of the environment that carry domain knowledge about which regions of the visual input are important for the downstream task. Such representations are crucial in RL because they enable agents to tractably deal with high-dimensional image observation spaces.
Prior works have shown that self-supervised learning with data augmentation helps achieve good performance in image-based RL. Contrastive Unsupervised RL (CURL) [3] employs a contrastive learning objective as an auxiliary loss to learn representations for off-policy RL. RL with Augmented Data (RAD) [24] and Data Regularized Q-Learning (DrQ) [4] use simple image augmentations such as random cropping and color jittering to provide strong regularization for learning representations invariant to visual perturbations. ViSaRL does not use data augmentation directly in the value function or policy update. Instead, saliency augmentation is introduced during the visual encoder pretraining phase.
Nair et al. [25] and Karamcheti et al. [26] propose to combine internet scale language and vision datasets with different self-supervised objectives to learn visual representations generally applicable across all robot tasks. While they focus on learning general visual representations, ViSaRL augments small task-specific datasets with saliency information to improve pretrained visual representations.
@@
III VISUAL SALIENCY-GUIDED REINFORCEMENT LEARNING
@@
Figure 3: ViSaRL Approach. We pretrain a MultiMAE [28] Transformer model on a dataset of paired RGB images and their saliency maps. MultiMAE employs a self-supervised training objective in which masked patches for both input modalities are reconstructed given only the visible patches. The pretrained model is frozen and used for extracting latent image representations during task learning. There is no input masking during downstream RL.
We propose ViSaRL, a simple approach for incorporating human-annotated saliency to learn representations for visual control tasks. ViSaRL can be implemented on top of any standard RL algorithm for learning a policy. It aims to learn representations that encode useful task-specific inductive biases from human saliency maps. ViSaRL consists of three learned components: a saliency predictor $g_{\phi}$, an image encoder $f_{\theta}$, and a policy network $\pi_{\psi}$. Figure 1 provides an overview of the ViSaRL architecture. We will elaborate on each component in the following subsections.
@@
Pretraining Visual Representation. We use our trained $g_{\phi}$ to pseudo-label an offline image dataset collected using any behavior policy (random, replay buffer, expert demonstrations, etc.) with saliency information. We then use the paired image and saliency dataset to pretrain an image encoder, $f_{\theta}$, for computing image embeddings. We experiment with two models for our backbone visual encoder, CNN and Transformer encoder, and investigate different techniques for augmenting each with saliency input. To add saliency to a CNN architecture, we can use saliency as a soft continuous mask or simply add saliency as an additional channel per pixel. For a Transformer encoder, we propose to pretrain the model with saliency as an additional input using a masked reconstruction objective.
Masked autoencoders (MAE) [30] are an effective and scalable approach for learning visual representations. MAE masks out random patches of an image and reconstructs the masked patches using a Vision Transformer (ViT) [31]. An image $I\in\mathbb{R}^{H\times W\times C}$ is processed into a sequence of 2D patches $h\in\mathbb{R}^{K\times(P^{2}C)}$ where $P$ is the patch size and $K=HW/P^{2}$ is the number of patches. A subset of these patches are randomly masked out with a masking ratio of $m$. Only the visible, unmasked patches are used as input to the ViT encoder. Masking reduces the input sequence length and encourages learning global, contextualized representations from limited visible patches.
The image patches are first embedded via a linear projection and added to positional embeddings. The resulting tokens are processed via a series of Transformer blocks. Finally, a ViT decoder reconstructs the original input by processing all of the tokens including the encoded visible patches and placeholder mask tokens. Following He et al. [30], we employ a high masking ratio $m=0.75$ and a heavy-encoder, light-decoder architecture to enable efficiently learning good representations.
@@
8:     Select action $a=\pi_{\psi}(f_{\theta}(o,g_{\phi}(o)))$
9:     Optimize $\mathcal{L}_{RL}$ wrt to $\psi$
10:end for
@@
Downstream Policy Learning. After pretraining the MultiMAE model, we freeze the encoder and use it to compute latent representations of environment observations for policy training. ViSaRL is not only compatible with online RL algorithms such as Soft-Actor Critic (SAC) [32] in which the agent learns through environment interactions but also imitation learning from expert demonstrations. Image inputs are not masked during policy learning. We average the patch embeddings to generate a global image representation. We also tried using the global learned token embedding, similar to a CLS token in ViT, which yielded similar results. The full procedure for ViSaRL is summarized in Algorithm 1.
@@ IV EXPERIMENT SETUP
To demonstrate the effectiveness of using human-annotated saliency information to enhance visual representations for task learning, we show quantitative results of our approach with two different encoder backbones, CNN and Transformer, across multiple simulated environments including the Meta-World manipulation [6] and DMC benchmarks [21] and real-robot manipulation with a Kinova Jaco 2 arm. We train the downstream policy using SAC [32] for the simulation experiments and behavioral cloning with expert demonstrations for the real robot experiments.
Saliency Map Annotation. We created a simple web-based user interface to collect saliency annotations shown in Figure 2. An annotator clicks on the pixels in the image that they think are relevant for performing the given downstream task. The interface creates a Gaussian centered around the clicked pixel with $\sigma=10$ on an input image of resolution $256\times 256\times 3$. For manipulation tasks, the salient regions could include the robot end-effector position, task specific objects, and the goal location.
@@
DMControl RGB (CNN) CURL RAD RGB(S) RGB+Saliency (PO) RGB+Saliency (Ours)
Walker Walk $50$±$\pm 25$ $895$±$\pm 45$ $912$±$\pm 38$ $914$±$\pm 32$ $918$±$\pm 36$ $\boldsymbol{925}$±$\pm 24$
Cartpole Swing $440$±$\pm 43$ $820$±$\pm 62$ $850$±$\pm 52$ $838$±$\pm 42$ $880$±$\pm 37$ $\boldsymbol{914}$±$\pm 16$
Ball Catch $325$±$\pm 55$ $944$±$\pm 33$ $932$±$\pm 52$ $939$±$\pm 58$ $946$±$\pm 62$ $\boldsymbol{962}$±$\pm 14$
Finger Spin $200$±$\pm 180$ $910$±$\pm 50$ $922$±$\pm 32$ $933$±$\pm 38$ $936$±$\pm 16$ $\boldsymbol{947}$±$\pm 52$
Average $254$±$\pm 76$ $892$±$\pm 48$ $904$±$\pm 44$ $906$±$\pm 42$ $920$±$\pm 38$ $\boldsymbol{937}$±$\pm 26$
TABLE II: Average evaluation return on four DMC Benchmark tasks. Average over 5 seeds each.
Quality of Saliency Prediction Model. We find that zero-shot evaluation using a pretrained PiCANet and even state-of-the-art Vision Language Models fail to correctly identify the task-relevant regions of the image necessitating the use of human-annotation to guide the saliency training. Thus, we train a new saliency model for each task. However, we emphasize that we can learn a saliency predictor that generalizes reasonably to unseen environment states with just five annotated examples (Figure 5). We leave training a multi-task saliency predictor for future work.
Figure 5: Saliency predictions on unseen environment states. The first row is the raw RGB observations. The second row is the predicted saliency and the last row is the RGB image masked with the predicted saliency.
We apply random vertical and horizontal flip for data augmentation to prevent overfitting on our small dataset. We measure the quality of the saliency model on two evaluation metrics following prior work [29]: F-measure score ($F_{\zeta}$) and Mean Absolute Error (MAE).
$F_{\zeta}=\frac{(1+\zeta^{2})Precision\times Recall}{\zeta^{2}Precision+Recall},$ (1)
where $\zeta^{2}$ is set to 0.3 [29]. Saliency maps are binarized before computing $F_{\zeta}$. MAE computes the average absolute per-pixel difference between the predicted saliency maps and ground truth saliency maps. On our test dataset, we obtain an $F_{\zeta}$ score of $0.78\pm 0.02$ and MAE of $0.004\pm 0.003$ averaged across all tasks, consistent with results reported in prior work [29]. Qualitative examples of the predicted saliency maps for the real-robot experiments are shown in Figure 5; further analysis can be found on the project page.  [^1] The code implementation for reproducing the results and additional analysis can be found on the project website: https://sites.google.com/view/visarl.
@@ V SIMULATION EXPERIMENTS
Figure 4, Table I and II summarize our main findings on four Meta-World robot manipulation tasks and five DMC tasks (Figure 6) using CNN and Transformer encoder backbones. The downstream policy is trained using Soft-Actor Critic (SAC). We additionally compare against two state-of-the-art methods for visual representation learning: CURL [3], a contrastive representation learning method and RAD [24], a method that combines various image augmentations such as color jitter and random crop to induce visual invariances in the learned representation.
Saliency input improves downstream task success rates. Incorporating saliency improves the task success rate in Meta-World using CNN and Transformer encoders by 13% and 18% respectively over the next best baseline. For DMC environments, we observe a 256% relative improvement in average return when using saliency input versus without. Our Transformer encoder results in an average 4% relative improvement in environment returns across all tasks over the next best prior work with a 7.5% improvement in the Cartpole Swing task.
@@ V-A CNN Encoder
V-B MultiMAE Transformer
For MultiMAE experiments, the encoder is pretrained using an offline dataset of (image, saliency map) pairs. We use a 4-layer ViT encoder and a 3-layer ViT decoder with a patch size of $8\times 8$ pixels. We pretrain the model for 400 epochs on images collected from a random policy. The encoder weights are frozen during downstream reinforcement learning, decoupling the representation and policy learning. We compare MultiMAE representations pretrained with RGB only (RGB) and both RGB and saliency (RGB+Saliency (PO), RGB+Saliency (Ours)). RGB+Saliency (PO) uses saliency only during pretraining while RGB+Saliency (Ours) uses saliency in both pretraining and downstream RL.
Using saliency in both pretraining and inference yields the best performance. We compare the full ViSaRL method (RGB+Saliency (Ours)) to pretraining using only the RGB images (RGB) in Tables I and II demonstrating that multimodal pretraining with saliency information significantly outperforms single modality pretraining by at least a 10% margin across all tasks. Notably, RGB achieves only 19% success on Faucet Open, while our approach solves the task with 62% success rate. Using saliency as an input for both pretraining and downstream RL (RGB+Saliency (Ours)) improves task success rate over RGB+Saliency (PO) because there are new observations during online training that were not in the pretraining dataset and the encoder could benefit from the saliency input to generate a better state representation.
ViSaRL representations generalize to unseen environments. We evaluate the generalizability of our learned representations on the challenging random colors and video backgrounds benchmark from DMControl-GB [33]. In DMControl-GB, agents trained in the original environment are evaluated on their generalization to the same environment with visually perturbed backgrounds using randomized color and video overlays (see Figure 7). ViSaRL significantly outperforms the baselines across all tasks as shown in Table III, with an average 19% and 35% relative improvement respectively for the color and video settings. We observe that our learned saliency predictor is robust to the visual perturbations and thus the noisy background is filtered before generating the image embedding.
Human-annotated saliency improves performance compared to depth and surface normals. We conduct ablation experiments to compare saliency versus other mid-level input modalities such as depth and surface normals proposed by Sax et al. [27]. We substitute saliency with these other modalities as input to the MultiMAE. In Table IV, we observe that neither depth nor surface normal features alone improves task success over just using RGB image input. By contrast, adding saliency as an additional modality consistently improves task success suggesting that human-annotated saliency information can help learn better visual representations compared to other input modalities.
@@ V-B MultiMAE Transformer
TABLE III: Average return of ViSaRL and baseline methods on the color and video environments from DMC-GB.
@@
TABLE IV: Human-annotated saliency versus depth and surface normals (SN) as input modalities to MultiMAE model.
@@ VI REAL ROBOT EXPERIMENTS
TABLE V: Task success rates in real-world tabletop manipulation tasks for RGB and RGB+Saliency with MultiMAE.
ViSaRL scales to real-robot tasks and is robust to distractor objects. Videos of evaluation trajectories for each task can be found on the project website. In Table V, we report the task success rates on 10 evaluation rollouts. We compare MultiMAE Transformer representations trained using RGB and RGB+Saliency. Even on the easier task of Pick Apple, using saliency augmented representations improves the success rate. On tasks with distractor objects and longer-horizon tasks such as Put Apple in Bowl, saliency-augmented representations nearly double the task success rate.
@@ VII CONCLUSION
In this paper, we only considered static frame saliency maps for single-object manipulation tasks. We plan to extend our approach to handle longer-horizon multi-object tasks using video saliency models [37] which can learn to encode more flexible temporal saliency representations across a sequence of frames. This extension could be implemented by asking the human users to watch video clips of the trajectories and annotate saliency over these clips. Another interesting direction is to investigate augmenting large-scale vision datasets with saliency to see whether saliency can benefit visual representations more generally.
APPENDIX
Saliency predictor training details. Since we use a small dataset for training the saliency predictor, we apply data augmentation to prevent overfitting and improve robustness of the model. We apply VerticalFlip, HorizontalFlip and ColorJitter(brightness=0.4, contrast=0.4, saturation=0.4, hue=0.2). We use AdamW [38] optimizer with learning rate of 3e-4 and cosine learning rate annealing.
Implementation details. Table VI summarize the hyperparameters in ViSaRL and other baselines. ViSaRL is agnostic to the choice of downstream RL algorithm. For the CNN encoder experiments, we follow the implementation from [24, 21] and train both the encoder and policy end-to-end using Soft-Actor Critic. The encoder consist of a stack of 11 convolutional layers, each with 32 filters of 3 $\times$ 3 kernels, no padding, stride of 2 for the first and 1 for all others. This results in a feature map of dimension $32\times 12\times 12$ from an input image of shape $64\times 64\times 3$.
The policy head $\pi_{\theta}$ and action-value functions $Q_{\phi_{i}}$ are parameterized by multi-layer perceptrons. The policy head is composed of a linear projection of dimension 100 with normalization followed by 3 linear layers with 1024 hidden units each and a final linear output layer for the action prediction. As the action spaces of DMControl suite and Meta-World are continuous, the policy outputs the mean and variance of a Gaussian distribution over actions.
Hyperparameter Value
Action repeat 1
Discount factor $\gamma$ 0.99
Episode length 500
Replay buffer size 1M
Policy learning rate Adam(lr=1e-4, $\beta_{1}=0.9,\beta_{2}=0.999$)
$\alpha$ learning rate Adam(lr=3e-4, $\beta_{1}=0.5,\beta_{2}=0.999$)
Batch size 128
Target network update 2
Target network momentum $\tau$ 0.05
Environment steps 1M
TABLE VI: Hyperparameters used in SAC training.
Meta-World: The observation space is a $64\times 64\times 3$ image. The action space $\mathcal{A}\subset\mathbb{R}^{4}$, is the $\Delta(xyz)$ of the end-effector, and a continuous scalar value for gripper torque. Object and goal positions are randomized at the start of every episode to prevent exploiting spurious correlation
Figure 7: DMControl Generalization Benchmark (Top) Four continuous control tasks from the DMControl suite. (Middle and Bottom) We evaluate on the color_easy and video_easy settings to test the generalizability of the learned representations.
Figure 8: Robot Policy Architecture We use an LSTM policy with MLPs to embed the input and generate continuous actions. The input is the cropped RGB image from an external camera and the predicted saliency map. The visual input is contatenated with the proprioceptive information.
Real Robot Imitation Learning. The downstream policy is trained using standard imitation learning. We train the saliency predictor and Transformer encoder using the same procedure outlined for the simulation experiments. Given our pretrained visual encoder, we form the state representation as a concatenation of the visual embedding and the robot proprioceptive information (e.g. joint positions). This yields a 271-dimensional state representation.
The policy is implemented as an LSTM with 256-dimensional hidden states which autoregressively predicts the actions for the next $H$ timesteps where $H$ is a fixed window. The state is processed by a 2-layer input MLP with hidden sizes [256, 256]. The final hidden state of the LSTM is processed by a 2-layer MLP to predict a continuous action. The full model architecture is shown in Figure 8. The action space, $\mathcal{A}\in\mathbb{R}^{7}$, consists of $\Delta(x,y,z,\phi,\theta,\psi)$, and a continuous scalar value for gripper speed.

### v2

@@ I INTRODUCTION
A key ingredient in solving visual control tasks is to learn visual representations that capture useful features of the sensory input to simplify the decision-making process. Many works in the deep reinforcement learning (RL) community have proposed to learn such representations through various self-supervised objectives including contrastive learning [3] and data augmentation [4]. By contrast, we focus on self-supervision using saliency as additional human domain knowledge to inform the representation of task-relevant features in the visual input while filtering out perceptual noise.
@@
Figure 1: ViSaRL trains a saliency prediction model from a few human-annotated saliency maps. This model is used to augment an offline image dataset with saliency. A visual encoder is pretrained with the dataset and used during downstream policy learning to generate latent representations of the agent’s observations.
@@ II RELATED WORK
User Interfaces for Human Saliency. ViSaRL needs a small number of human-annotated saliency maps to bootstrap the saliency prediction network. Prior work used superpixel segmentation [22] to first divide each image into segments, and then asked humans to click on the segments that are salient [17]. However, that method requires manually checking and combining the segments that belong to the same object before showing the images to annotators, burdening system designers. As an alternative, Boyd et al. [23] used interfaces where the annotators created binary masks by simply clicking on images. We employ a similar but simpler interface: an annotator clicks on the salient parts of the image, and a Gaussian kernel is applied around selected pixels to achieve smooth saliency maps shown in Figure 2.
Representation Learning for RL. Saliency maps are representations of the environment that carry domain knowledge about which regions of the visual input are important for the downstream task. Such representations are crucial in RL because they enable agents to tractably deal with high-dimensional image observation spaces.
Prior works have shown self-supervised learning with data augmentation helps achieve good performance in image-based RL. Contrastive Unsupervised RL (CURL) [3] employs a contrastive learning objective as an auxiliary loss to learn representations for off-policy RL. RL with Augmented Data (RAD) [4] use simple image augmentations such as random cropping and color jittering as regularization to learn representations invariant to visual perturbations. ViSaRL does not use data augmentation directly in the value function or policy update. Instead, saliency augmentation is introduced during the visual encoder pretraining phase.
Nair et al. [24] and Karamcheti et al. [25] propose to combine internet scale language and vision datasets to learn visual representations applicable across all robot tasks. While they focus on learning general visual representations, ViSaRL augments small task-specific datasets with saliency information to improve pretrained visual representations.
@@
III VISUAL SALIENCY-GUIDED RL
@@
Figure 3: ViSaRL. We pretrain a MultiMAE [27] Transformer on a dataset of paired images and saliency maps. MultiMAE employs a self-supervised objective in which masked patches for both input modalities are reconstructed given only the visible patches. The pretrained model is frozen and used for extracting representations during task learning. There is no input masking during downstream RL.
We propose ViSaRL, a simple approach for incorporating human-annotated saliency to learn representations for visual control tasks. ViSaRL can be implemented on top of any standard RL algorithm for learning a policy. It aims to learn representations that encode useful task-specific inductive biases from human saliency maps. ViSaRL consists of three learned components: a saliency predictor $g_{\phi}$, an image encoder $f_{\theta}$, and a policy network $\pi_{\psi}$ shown in Figure 1. We will elaborate on each component in the following sections.
@@
Pretraining Visual Representation. We use our trained $g_{\phi}$ to pseudo-label an offline image dataset collected using any behavior policy (random, replay buffer, expert demonstrations, etc.) with saliency maps. We then use the paired image and saliency dataset to pretrain an image encoder, $f_{\theta}$. We experiment with two models for our backbone visual encoder, CNN and Transformer, and investigate different techniques for augmenting each with saliency input. To add saliency to a CNN, we can use saliency as a continuous mask or simply add it as an additional channel per pixel. For a Transformer encoder, we pretrain the model with saliency as an additional input using a masked reconstruction objective.
Masked autoencoders (MAE) [29] are an effective and scalable approach for learning visual representations. MAE masks out random patches of an image and reconstructs the masked patches using a Vision Transformer (ViT) [30]. An image $I\in\mathbb{R}^{H\times W\times C}$ is processed into a sequence of 2D patches $h\in\mathbb{R}^{K\times(P^{2}C)}$ where $P$ is the patch size and $K=HW/P^{2}$ is the number of patches. A subset of these patches are randomly masked out with a masking ratio of $m$. Only the visible, unmasked patches are used as input to the ViT encoder. Masking reduces the input sequence length and encourages learning global, contextualized representations.
The image patches are embedded via a linear projection and added to positional embeddings. The resulting tokens are processed via a series of Transformers. Finally, a ViT decoder reconstructs the original input by processing all of the tokens including the encoded visible patches and placeholder mask tokens. Following He et al. [29], we set a high masking ratio $m\!=\!0.75$ and a heavy-encoder, light-decoder architecture.
@@
8:     Select action $a\sim\pi_{\psi}(f_{\theta}(o,g_{\phi}(o)))$
9:     Optimize $\mathcal{L}_{RL}$ with respect to $\psi$
@@
Downstream Policy Learning. After pretraining the MultiMAE model, we freeze the encoder and use it to compute latent representations of environment observations for policy training. ViSaRL is not only compatible with online RL algorithms such as Soft-Actor Critic (SAC) [31] in which the agent learns through environment interactions but also imitation learning from expert demonstrations. Image inputs are not masked during policy learning. We average the patch embeddings to generate a global image representation. The full procedure for ViSaRL is summarized in Algorithm 1.
@@ IV EXPERIMENT SETUP
To demonstrate the effectiveness of using human-annotated saliency information to enhance visual representations for task learning, we show quantitative results of our approach with two different encoder backbones, CNN and Transformer, across multiple simulated environments including the Meta-World manipulation [6] and DMC benchmarks [21] and real-robot manipulation with a Kinova Jaco 2 arm. We train the downstream policy using SAC [31] for the simulation experiments and behavioral cloning with expert demonstrations for the real robot experiments.  [^1] The code implementation for reproducing the results and additional analysis can be found on: https://liralab.usc.edu/visarl/.
Saliency Map Annotation. We created a simple user interface to collect saliency annotations shown in Figure 2. An annotator clicks on the pixels in the image that they think are relevant for performing the given downstream task. The interface creates a Gaussian centered around the clicked pixel with $\sigma=10$ on an input image of resolution $256\times 256\times 3$.
@@ V SIMULATION EXPERIMENTS
Figure 4 and Table I summarize our main findings on 4 Meta-World robot manipulation tasks and 5 DMC tasks using CNN and Transformer backbones. We compare against two state-of-the-art methods for visual representation learning: CURL [3], a contrastive representation learning method and RAD [4], a method to combine various image augmentations to induce visual invariances in the learned representations.
Saliency input improves downstream task success rates. Incorporating saliency improves the task success rate in Meta-World using CNN and Transformer encoders by 13% and 18% respectively over the next best baseline. For DMC environments, we observe a 256% relative improvement in average return when using saliency input. Our Transformer encoder results in an average 4% relative improvement in environment returns across all tasks over the next best baseline with a 7.5% improvement in Cartpole Swing.
@@ V-A CNN Encoder
TABLE II: Average return of ViSaRL and baseline methods on the color and video environments from DMC-GB.
V-B MultiMAE Transformer
We compare MultiMAE representations pretrained with RGB only (RGB) and both RGB and saliency (RGB+Saliency (PO), RGB+Saliency (Ours)). RGB+Saliency (PO) uses saliency only during pretraining while RGB+Saliency (Ours) uses saliency in both pretraining and downstream RL.
Using saliency in both pretraining and inference yields the best performance. We compare the full ViSaRL method (RGB+Saliency (Ours)) to pretraining using only the RGB images (RGB) in Tables I demonstrating that multimodal pretraining with saliency information significantly outperforms single modality pretraining by at least a 10% margin across all tasks. Notably, RGB achieves only 19% success on Faucet Open, while our approach solves the task with 62% success rate. Using saliency as an input for both pretraining and downstream RL (RGB+Saliency (Ours)) improves task success rate over RGB+Saliency (PO) because there are new observations during online training that were not in the pretraining dataset.
ViSaRL representations generalize to unseen environments. We evaluate the generalizability of our learned representations on the challenging random colors and video backgrounds benchmark from DMControl-GB [32]. In DMControl-GB, agents trained in the original environment are evaluated on their generalization to the same environment with visually perturbed backgrounds using randomized color and video overlays. ViSaRL significantly outperforms the baselines across all tasks as shown in Table II, with an average 19% and 35% relative improvement respectively for the color and video settings.
Human-annotated saliency improves performance compared to depth and surface normals. We conduct ablation experiments to compare saliency versus other mid-level input modalities such as depth and surface normals proposed by Sax et al. [26]. We substitute saliency with these other modalities as input to the MultiMAE. In Table III, we observe that neither depth nor surface normal features alone improves task success over just using RGB image input. By contrast, adding saliency as an additional modality consistently improves task success suggesting that human-annotated saliency information can help learn better visual representations compared to other input modalities.
@@ V-B MultiMAE Transformer
TABLE III: Human-annotated saliency versus depth and surface normals (SN) as input modalities to MultiMAE model.
@@ VI REAL ROBOT EXPERIMENTS
TABLE IV: Task success rates in real-world tabletop manipulation tasks for RGB and RGB+Saliency with MultiMAE.
ViSaRL scales to real-robot tasks and is robust to distractor objects. Videos of evaluation trajectories for each task can be found on the project website. Table IV reports the task success rates on 10 evaluation rollouts. Even on the easier Pick Apple task, using saliency augmented representations, RGB+Saliency, improves the success rate over RGB. On tasks with distractor objects and longer-horizon tasks such as Put Apple in Bowl, saliency-augmented representations nearly double the success rate.
@@ VII CONCLUSION
In this paper, we only considered static frame saliency maps for single-object manipulation tasks. We plan to extend our approach to handle longer-horizon multi-object tasks using video saliency models [36] which can learn to encode more flexible temporal saliency representations across a sequence of frames. This extension could be implemented by asking the human users to watch video clips of the trajectories and annotate saliency over these clips.